0.12x+0.7(x-2)=0.01(3x-4)
Trying to solve this in fraction form stuck at what to do? I know i need to get rid of the decimals..

Answer by fractalier(6550)

We can clear all the decimals by multiplying everything by 100...so from
0.12x+0.7(x-2)=0.01(3x-4)
we get
12x + 70(x-2) = 3x - 4
12x + 70x - 140 = 3x - 4
82x - 140 = 3x - 4
79x = 136
x = 136/79